Project

General

Profile

NDOS Commissioning Meeting 2011/10/24

Attendance: Jon, Martin, Peter, Leon, Susan, Nick, Raphael, Athans, Enrique,

Agenda:

  1. APD Redistribution Status & Update
  2. Beam & Detector Status
  3. DAQ/Electronics Status & Plans
  4. DCS Status & Plans
  5. AOB

APD Redistribution Status & Update: Denis and Mat will continue the APD redistribution starting tomorrow and expect to finish by the end of the week.

Beam & Detector Update

  • cooling and FEB drive currents were in alarm this morning, it would be good to know why/what happened.
  • Martin will look into it

DAQ/Electronics Status & Plans

  • Leon: none that he is aware of
  • Peter: perhaps some tests later this week, but nothing definitive

DCS Status & Plans

  • no plans

AOB

  • none